Part A What must the initial speed of a lead bullet be at a temperature of 23.0 ∘C so that the heat developed when it is brought to rest will be just sufficient to melt it? Assume that all the initial mechanical energy of the bullet is converted to heat and that no heat flows from the bullet to its surroundings. (Typical rifles have muzzle speeds that exceed the speed of sound in air, which is 346 m/s at 23.0 ∘C ). v = m/s
Answer #1

given that all the mechanical energy is converted to heat

kinetic energy = heat energy

1/2*m*v^2 = m*c*\Delta t

melting temperature of lead = 327.5 celsius , c = 128

hence

v = 279.19 m/s

1/2*v^2 = c*\Delta t